I need to know how this answer derived: I figured it out, but I do not know how I figured it out.
76=2(11) + 2(11) +5
The length of a rectangle is 5 in more than twice the width. If the perimeter of the rectangle is 76 in, what are the dimensions of the rectangle?

Answer by oberobic(2304)

L = 2W + 5
.
P = 2(L+W) = 76
L+W = 38
so
L = 38-W
.
L = L
so
2W + 5 = 38-W
.
3W = 33
.
W = 11
.
L = 2W + 5
L = 2(11) + 5
L = 27
.
Check the perimeter
P = 2(27+11) = 2*38 = 76
Correct.
.
Answer: The rectangle has length = 27 inches and width = 11 inches.
.
Done.